OpenStudy (sweetburger):

Is this how you would determine the overall charge on a compound? Alright this is what I did. I took NO3- and I used the formula (Valence Electrons-Bonding Electrons/2-Non Bonding Electrons) I got 2 of the Oxygen atoms having a -1 charge, the other oxygen having a 0 charge, and finally the Nitrogen having a +1 charge. So i added them up and its overall -1. Is this how you would normally determine the overall charge of a compound?

OpenStudy (abmon98):

It Depends if it is a neutral compound then the sum of oxidation states must equal zero. and if its a compound with charge either + or - the sum of oxiddation states of atom must equal to charge of compund.

OpenStudy (somy):

How did you get N charge in NO3- to be +1?

OpenStudy (somy):

Oxygen atom has a fixed oxidation state of -2

OpenStudy (abmon98):

Oxygen has a charge of -2, 3 oxygen atoms are present -2*3=-6 Charge on Nitrate ion is -1 therefore Nitrogen must have an oxidation of +5

OpenStudy (somy):

NO3- means overall charge is -1 for O its fixed to have -2 as its charge thus X+ (-2*3) = -1 X -6 = -1 X= -1+6 X= +5

OpenStudy (somy):

overall charge of a compound depends on 

OpenStudy (somy):

if there is -1, -2 , -whatever or +1, +2 or +whatever that is going to be the overall charge of a compound if there is no such thing, then overall charge of a compound would always be 0

OpenStudy (sweetburger):

Yes, @Somy I understand that I wanted to determine how one would see why the compound has a negative or positive charge. Thank you for showing me that it just has to do with oxidization numbers. Btw I got Nitrogen was +1 using the formula I provided...

OpenStudy (somy):

idk where you got that formula from but this is how we usually do it @sweetburger :)

OpenStudy (sweetburger):

Organic Chemistry book by John McMurray I think it may just deal with reasonance I was trying to see if it had any other uses

OpenStudy (somy):

hmm i see well it's not fitted here i guess